OpenStudy (anonymous):

The girls group rented 2 canoes and 3 inner tubes for a total cost of $80. The boys group rented 1 canoe and 7 inner tubes for a total cost of $95. What is the cost to rent 1 canoe? A. $18 B. $15 C. $25 D. $10 <33 please help

OpenStudy (anonymous):

1) 2a+3b = 80 2) a+7b = 95 rewrite 2) a = 95-7b plugin into 1) 190-14b+3b = 80 11 b = 110 canoe must be 25 bucks if the inner tubes are 10

OpenStudy (lilg132):

x = canooes y = inner tubes 80 = 2x + 3y 95 = x + 7y (multiply by 2) 80 = 2x + 3y 190 = 2x + 14y minus from eachother 110 = 11y 110 / 11 = 11y / 11 10 = y inner tubes cost $10 each now we can work out how much the canooes cost 80 = 2x + 3y we replace y with the amount we found for how much the inner tubes cost which is 10 80 = 2x + 3(10) which is 80 = 2x + 30 which now turns to 80 - 30 = 2x 50 = 2x 50 / 2 = 2x / 2 25 = x now we know the price for the canooes is $25 each and to check we can use the amounts in the second equation 95 = x + 7y replace x with the price of the canooes and replace y with the price of the inner tubes 95 = 25 + 7(10) 95 = 25 + 70 and we can see the figures are right because they equal to 95 remember to click good answer